OpenStudy (toxicsugar22):

The cost in thousand of dollers for extracting p% of a precius ore from a mine is given by this equation C(P)=360P/100-P What input values make sence for C? If 15=p then c is if 25=p then c is if 40= p then c is if 50 =p then c is

myininaya (myininaya):

i don't see what comes after then c is

OpenStudy (toxicsugar22):

my question is what is c if p is 15, 25 and so on

myininaya (myininaya):

oh i see

myininaya (myininaya):

if P=15, then C(15)=360(15)/100-15 =36(15)/10-15 =36(3)/2-15 =18(3)-15 =54-15 =39

myininaya (myininaya):

you do the same for the other 3 values for P just plug them in

OpenStudy (toxicsugar22):

i dont thats how to do it

OpenStudy (toxicsugar22):

no im sorry

OpenStudy (toxicsugar22):

The cost in thousand of dollers for extracting p% of a precius ore from a mine is given by this equation C(P)=360P/100-P complete the table showing the cost of extracting various percents of the ore. If 15=p then c is if 25=p then c is if 40= p then c is if 50 =p then c is

myininaya (myininaya):

you see P=25, take this 25 and plug it wherver you see a P in C(P)=360P/100-P

OpenStudy (toxicsugar22):

ok i did that

OpenStudy (toxicsugar22):

thankss